OpenStudy (anonymous):

Solve 16^log4 x = 25

OpenStudy (agent0smith):

\[\huge 16^{\log_4 x} = 25\]

OpenStudy (hitaro9):

Is that 16 to the log of 4x, or 16 to log base 4 of x?

OpenStudy (isaiah.feynman):

@milkacha Please we'd like to be sure.

OpenStudy (agent0smith):

In either case, you could start by taking logs of both sides.

OpenStudy (shamil98):

could you screenshot it or take a picture?..

OpenStudy (isaiah.feynman):

Please attach a picture of the original question.

OpenStudy (hitaro9):

Where are you getting the problem from? Is it online written exactly like that, is it from a book?

OpenStudy (hitaro9):

We basically need to know where the 4 is relative to the x

OpenStudy (hitaro9):

If the 4 is a little bit lower than the x, that's a different problem from if the 4 is right next to the x.

OpenStudy (agent0smith):

Take logs \[\huge \log 16^{\log_4 x} =\log 25\]use log laws\[\huge \log_4 x \log 16 =\log 25\]divide both sides by log16\[\huge \log_4 x =\frac{ \log 25 }{ \log 16 }\] and I'll stop here till you clarify the problem. The solution is the same till here in either case.

OpenStudy (agent0smith):

You still have to put both sides to the power of 4, to find x.

OpenStudy (anonymous):

Yep all good I can take it from here, thanks alot.